# Partial domain adaptation enables cross domain cell type annotation between scRNA-seq and snRNA-seq

### v2

@@ 5 Conclusion
In this work, we introduced ScNucAdapt, a novel framework designed for cross-domain cell-type annotation between scRNA-seq and snRNA-seq datasets. To the best of our knowledge, ScNucAdapt is a method to explicitly address this challenge by integrating principles from partial domain adaptation with dynamic clustering and distribution alignment through Cauchy–Schwarz divergence. This design enables ScNucAdapt to selectively transfer relevant information from annotated data to unlabeled data, mitigating the impact of non-overlapping cell types and modality-specific biases.
Comprehensive experiments across multiple tissues and conditions—including bladder, kidney, tumor, and cortical datasets. Demonstrate that ScNucAdapt consistently outperforms existing annotation tools such as SingleCellNet and ScMap, as well as domain adaptation methods ScAdapt. The ablation and sensitivity analyses further confirm that both the CS divergence and dynamic cluster selection modules are essential for achieving robust generalization and stability across heterogeneous datasets.
Overall, ScNucAdapt provides a practical and scalable solution for leveraging well-characterized scRNA-seq data to annotate more challenging snRNA-seq datasets, particularly from frozen or difficult-to-dissociate tissues. Future extensions may focus on addressing label noise in source data, domain adaptation overfitting, novel cell type discoveries, and class imbalancing in each domain.
@@
No competing interest is declared.
@@ Acknowledgments
The authors thank the anonymous reviewers for their valuable suggestions. This work is supported in part by funds from the National Natural Science Foundation of China(62131004, 62531002, 62306051, 62481540175, 62276035), the Taishan Scholars Foundation of Shandong Province(tsqn202507225), and the Natural Science Foundation of Chongqing (CSTB2025NSCQ-GPX0857). The Fundamental Research and the Scientific and the Technological Research Program of Chongqing Municipal Education Commission (KJQN202300718).

### v3

@@ 5 Conclusion
In this study, we introduced ScNucAdapt, a novel cross-domain annotation framework designed specifically for transferring cell type labels between paired or unpaired scRNA-seq and snRNA-seq datasets. To the best of our knowledge, this is the first method to address the unique challenges of cross-annotation across these two sequencing protocols. ScNucAdapt tackles both distributional differences and label space mismatches between source and target domains through three key components: a shared encoder that projects both datasets into a common latent space, a dynamic clustering mechanism that adaptively identifies the unknown number of cell types in the target data via split-merge operations, and a Cauchy-Schwarz divergence-based matching strategy that aligns source classes with target clusters while minimizing negative transfer from non-shared cell types.
Extensive experiments on eight cross-domain annotation tasks spanning bladder, kidney, tumor, and mouse cortical tissues demonstrated that ScNucAdapt consistently outperforms existing methods, including scRNA-seq classifiers and domain adaptation baselines, in both accuracy and macro F1-score. The framework proves effective under both closed-set and partial-set scenarios, maintaining robust performance even when target label spaces are subsets of the source. Ablation studies confirmed the necessity of each proposed component, while sensitivity analyses showed that ScNucAdapt is robust to hyperparameter choices, requiring minimal tuning in practice.
Our scalability analysis on simulated datasets confirmed that ScNucAdapt exhibits linear memory scaling and manageable runtime for datasets up to 16,000 cells, with the primary computational bottleneck being the GMM clustering and split-merge operations performed each epoch. For larger datasets, we recommend reducing the frequency of these operations or exploring approximate clustering variants.
Despite these strengths, several promising directions remain for future work. These include handling label noise in source annotations, extending the framework to discover novel cell types in target domains through open-set or universal domain adaptation, addressing heterogeneous feature spaces where gene sets differ substantially between datasets, mitigating overfitting in high-dimensional sparse spaces, and developing strategies to better handle imbalanced cell type distributions. Addressing these challenges will further enhance the applicability and robustness of cross-domain annotation methods in real-world single-cell and single-nucleus studies.
In summary, ScNucAdapt provides a powerful and flexible solution for integrating and annotating scRNA-seq and snRNA-seq data, enabling more consistent and reliable interpretation of cellular identities across experimental protocols and tissue conditions.
@@
The authors declare that there are no competing interests.
@@ Acknowledgments
This work is supported in part by funds from the National Natural Science Foundation of China (62131004, 62531002, 62306051, 62481540175, 62276035), the Taishan Scholars Foundation of Shandong Province (tsqn202507225), and the Natural Science Foundation of Chongqing (CSTB2025NSCQ-GPX0857). The Fundamental Research and the Scientific and the Technological Research Program of Chongqing Municipal Education Commission (KJQN202300718).